OpenStudy (anonymous):

Solve the formula P = 2l + 2w for w.

OpenStudy (anonymous):

First isolate your "w" value by moving 2l to the other side by subtracting it from both sides: P - 2l = 2w Now divide both sides by 2: P/2 - 2l/2 = 2w/2 which is the same as: P/2 - l = w So w = P/2-l